From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sun Mar 02 2003 - 07:52:12 EST
------------------------------------------------------------ Date: Sun Mar 2 07:52:12 2003, Shift: 00:00-08:00 Supervisor: BHS, with: Log book#14 pages 43 - 43 Shift Summary: ------------------------------------------------------------ We've had two stores this shift - one that was just starting when I came and lasted until 01:23, when RHIC developed a blue vacuum valve interlock. This took a few hours to sort out, and another store was achieved at 05:15. Still running on that one, having collected almost 300k trig 6 :-) Data runs: Setting Trig 3 Trig 6 8163 FS 12, 1/4A 50k MRS 40, 1050B 136k 8170 FS 12, 1/4A 55k MRS OFF due to high bg. 8171 FS 4, 1/4A MRS 40, 1050 B Still running IMPORTANT NOTE: I noticed when I came that run 8163 said FS 8deg in the DAQ - this is wrong. The correct value is 12deg. Then I, in my tiredness, forgot to cange this for the next run so 8170 also has the wrong value. Can someone please change runs 8163 and 8170 to 12deg runs in the db? (Run 8171 is correctly labeled a 4deg run...) Also, the mainframe for TOFWII was turned on by ZY on the last shift, but was off again when I looked at it so clearlty it's unstable. I have not tried to play around with it to keep it on, but leave this for the experts. The printer up here has taken leave of the net. It says "online" and IDLE, but I can't ping it or otherwise reach it over the net (yes, I've tried restarting and yanking the cables ;-) So I haven't been able to print any supermon plots etc. tonight - look at them on the web instead if you are interested. ...plus the odd T5 trip, hopefully quickly detected and fixed. ------------------------------------------------------------
This archive was generated by hypermail 2.1.5 : Sun Mar 02 2003 - 07:56:37 EST